how many pounds of candy that sells for 80 cents a pound could be mixed with candy that sell for 1.50 a pound to make a 20 pounds of a mixture to sell at 1.01 a pound
Answer by robertb(5830)

0.80x + (20-x)1.50 = 1.01*20
==> 0.80x + 30 - 1.50x = 20.2
==> -0.7x = -9.8
==> x = 14 pounds of the 80 cent per pound candy, and
20 - x = 6 pounds of the 1.50 $ per pound candy.